# Automatic Album Sequencing

### v1

@@ 1 Introduction
Album sequencing is the process of taking a music album and ordering it so that listening to it in that order produces a desired emotional response in the listener. Despite its importance in producing an impactful music album, album sequencing has received comparatively little attention from the artificial intelligence community.
@@ 2 User Interface
Our user interface is shown in Figure 1. This interface is built in Python [^1] https://www.python.org/ using the streamlit library [^2] https://streamlit.io/ and thus requires comparatively minimal effort to use. The interface allows a user to upload files, produce several fittings, and then visualize the results of these fittings.

### v2

@@ 1 Introduction
Playlist sequencing is the process of taking a collection of music tracks and ordering them so that listening to them in that order produces a desired emotional response in the listener. When the collection of music tracks together form an album, this process is known as album sequencing. Despite its importance in producing impactful music albums and playlists [1, 2], this sequencing process has received little attention from the artificial intelligence community.
@@
2 Related Work
While there is a considerable body of work looking at music playlist continuation (e.g., [10, 11, 12, 13, 14]), work on album or playlist sequencing is more sparse. Even sparser is work that uses deep neural networks to accomplish this. These gaps persist despite several works showing the sequencing process’s importance and complex nature (see [1, 2, 15]) and an abundance of work on using deep learning for music recommendation applications (e.g., [16]). Most of the existing methods that perform playlist sequencing focus either on controlling the similarity of adjacent songs (e.g., [17, 18, 19, 20]) or on directly maximizing the user preference (e.g., [21]). Altogether, this leaves the most closely related works to ours being our previous work in this area, where we extracted a low-dimensional representation of songs and fit these representations to a set of learned template curves [3], and the concurrent work of Neto et al., who looked at patterns in song-sequencing using pre-defined high-level features [22].
@@ 3 User Interface
Our first contribution is the web-based user interface shown in Figure 1. This interface is built in Python [^1] https://www.python.org/ using the streamlit library [^2] https://streamlit.io/ and thus requires comparatively minimal effort to use. The interface allows a user to upload files, produce several fittings, and then visualize the results of these fittings.